The coil of wire in a galvanometer has a resistance of RC = 58.7Ω. The galvanometer exhibits a full-scale deflection when the current through it is 0.226 mA. A resistor is connected in series with this combination so as to produce a voltmeter. The voltmeter is to have a full-scale deflection when it measures a potential difference of 12.8 V. What is the resistance of this resistor?


1
Expert's answer
2021-03-02T18:03:56-0500

"I=\frac{12.8}{R+R_C}\to IR+IR_C=12.8\to R=\frac{12.8-IR_C}{I}="


"=\frac{12.8-0.000226\cdot 58.7}{0.000226}=56578\ (\Omega)" . Answer